OpenStudy (anonymous):

Which is an equation of a circle with center (-5, -7) that passes through the point (0, 0)? (x + 5)2 + (y + 7)2 = 74 (x - 5)2 + (y - 7)2 = 74 (x - 5)2 + (y - 7)2 = 37 (x + 5)2 + (y + 7)2 = 37

OpenStudy (anonymous):

(x-k)^+(y-h)^2=r2

OpenStudy (anonymous):

k is x and h is y and those are the centers

OpenStudy (mathmale):

First, would you look up (if necessary) and type out here the standard equation of a circle with center at (h,k) and radius r? joshfers is on the right track, but is not fully correct. Note that I'm addressing you, @fack44.

OpenStudy (anonymous):

(x - h)2 + (y-k)2 = r2

OpenStudy (anonymous):

right.I forgot the square on the first equation and put inverse the letters.Now just plug the centers that they gave you in the equation.

OpenStudy (mathmale):

sorry to be picky, but the SQUARE of (x-h) needs to be written properly: (x-h)^2 or \[(x-h)^2\] .... but not as (x-h)2. Or you could use the Draw utility (below):

OpenStudy (mathmale):

Please write or type the standard equation of a circle with center at (h,k) and radius r.

OpenStudy (anonymous):

so is it (x - 5)^2 + (y - 7)^2= 37

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I was hoping you'd write the standard equation of a circle with center at (h,k) and radius r:\[(x-h)^2+(y-k)^2 = r^2\]

OpenStudy (mathmale):

The center of your circle is (-5, -7). Thus, your h = -5 and your y = -7. substitute these into the standard equation (above).